OpenStudy (anonymous):

The EXACT VALUE of tan(5pi/12) is:

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@Homeworksucks

OpenStudy (usukidoll):

easier to convert into degrees using \[\frac{180}{\pi}\]

OpenStudy (usukidoll):

\[\frac{5 \pi}{12} \times \frac{180}{\pi}\] the pi's cancel out so in degree mode it's 75

OpenStudy (anonymous):

its multiple choice: a) 2+rad3 b) 2-rad3

OpenStudy (usukidoll):

ack.

OpenStudy (anonymous):

so....

OpenStudy (usukidoll):

when I took tan(75) I got 3.73 since radical 3 is already 1.73 you add 2.00 and it's 3.73.

OpenStudy (anonymous):

but thats not exact value doesnt exact value have radicals in the answer?

OpenStudy (usukidoll):

this stuff is radian mode... I don't know why though... degrees are better. based off the multiple choice you had wither 2 + radical 3 or 2 - radical 3 I already memorized that radical 3 is around 1.73, so if you add 2.00 it is 3.73 which is the same as 5pi/12 for tangent

OpenStudy (usukidoll):

try adding it on the calculator and compare 2+ radical 3 and tangent(75) in degree mode it's the same. The reason why I didn't put exact is too many decimal places.

OpenStudy (usukidoll):

I'll screen shot it for ya.

OpenStudy (usukidoll):

same thing. Mathematics witchcraft BABY!!!!!!! :D

OpenStudy (usukidoll):

it's A . B doesn't make sense 2 - 1.73 = 0.27